OpenStudy (anonymous):

when 10^93-93is expressed as a single whole #, what is the sum of its digits?

OpenStudy (anonymous):

I think the answer is 826. 10^93=1000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000 10^93-93=999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999907 9*91=819+7=826.

OpenStudy (anonymous):

!0^93 is 1 followed by 93 zeros. Subtract 93 from that, and you have 91 nines, 1 zero and 1 seven. So your answer = 91*9 + 7 = 826.